OpenStudy (anonymous):

integrate (x+sinx)/(1+cosx) from 0 to pi/2

OpenStudy (dumbcow):

hmm im not seeing an easy way to do this. i would split it into 2 fractions and integrate separately x/(1+cos) ---> use integration by parts sin/(1+cos) --> use substitution

OpenStudy (anonymous):

Looks like it might involve this substitution: tan(x/2) = t So right away, you have x = 2arctan(t), as well as 1/2 sec²(x/2) dx = dt dx = 2 cos²(x/2) dt Making use of some trig identities, you'll find a few expressions you'll need: cosx = cos(2 (x/2)) = cos²(x/2) - sin²(x/2) sinx = sin(2 (x/2)) = 2 sin(x/2) cos(x/2) Use the initial sub to draw yourself a triangle that satisfies the equation tan(x/2) = t. You'll find yourself with a triangle that has sin(x/2) = t/(t² + 1) cos(x/2) = 1/(t² + 1) So, you get cosx = (1 - t²)/(t² + 1)² sinx = 2t/(t² + 1)² Returning to the differential substitution, you have dx = 2 [(1 - t²)/(t² + 1)²]² dt So the integral becomes Int[ (x + sinx)/(1 + cosx) dx ] = ... = Int[ (2arctan(t) + 2t/(t² + 1)²) / (1 + (1 - t²)/(t² + 1)²) * 2 [(1 - t²)/(t² + 1)²]² dt ] Hard to make out? Here's what it should look like: